OpenStudy (anonymous):

Permutations questions? Four in total.. This may be confusing, but I am going to try and write it out.. The "E"s are supposed to represent Greek E's. 5 E 2n+1 evaluate n=1 3 E 3n evaluate n=1 Type the formula for the number of permutations of n things taken r at a time. P(n,r)=? Type the formula for the number of circular permutations of n things. Pc(n,n) = ? Thanks y'all for explaining!

OpenStudy (anonymous):

Let's take this a little bit at a time. First of all, how to do summations. \[\Large \sum_{n=1}^{5}2n+1\] means "Summation from n=1 to n=5 of (2n+1)"

OpenStudy (anonymous):

What you do is you start at n=1 and you go up by 1 each time until you get to 5. So, n=1, then n=2, then n=3, then n=4, and finally n=5. At each n, you do the 2n+1, and you add it to the total.

OpenStudy (anonymous):

Any questions? Try it and let me see your attempt.

OpenStudy (anonymous):

Oh crud. Okay. Hmm. You have the first summation. Starting with 1, you substiute it in. 2(1)+1=3, 2(2)+1=5, 2(3)+1=7, 2(4)+1=9, 2(5)+1=11. Then you add them all? 3+5+7+9+11=35. So your answer is 35? Am I close..?

OpenStudy (anonymous):

So goooood! You're a boss.

OpenStudy (anonymous):

Try the next summation. \[\Large \sum_{n=1}^{3}3n\]

OpenStudy (anonymous):

Hahahaha. I've been told. Just kiddingg. ;) And okay. Let's go.. You will go from 1-3 because 3 is at the top. Subsituting starting with 1 is.. 3(1)=3, 3(2)=6, 3(3)=9. 3+6+9=18. Maybe? What does the n=1 at the bottom mean? What I have noticed is that is almost always 1.

OpenStudy (anonymous):

It does two things. It tells you which variable is changing each time, so in this case it's n. Sometimes it might say x=1 or i=1, which would mean that x or i was changing. The other thing is that it tells you which number to START at, so if it doesn't say n=1, but n=4, then you would start at 4 instead of 1.

OpenStudy (anonymous):

Ohh. Okay. That's probably a good thing to know :) Thanks!

OpenStudy (anonymous):

And yes, most examples will start at 1, but especially later on, they won't all start at 1. Let me show you an example to help you see one that ISN'T like that: \[\Large \sum_{i=3}^{6} 2i = 2(3) + 2(4) + 2(5) + 2(6)\] = 6 + 8 + 10 + 12 = 36

OpenStudy (anonymous):

Holy moly. Well, I'm glad I'm not at that yet.. Haha

OpenStudy (anonymous):

And thank you soo much!! :) I really appreciate it!

OpenStudy (anonymous):

Okay, next part. Permutations.

OpenStudy (anonymous):

Oh great. But no time like the present I suppose.

OpenStudy (anonymous):

I'd like to use an example. Imagine that I have a bag with 5 marbles in it, and they are 5 different colors. How many possible colors could I pull for the first marble?

OpenStudy (anonymous):

5 colors

OpenStudy (anonymous):

Cool. So 5 possibilities if I'm just drawing one marble. Now, imagine I've pulled one marble. How many possibilities are there for the second marble I draw?

OpenStudy (anonymous):

4 colors

OpenStudy (anonymous):

Exactly. So if I pull 2 marbles, there are 5 possibilities for the first marble and only 4 for the second marble. So the number of possible outcomes are 5*4

OpenStudy (anonymous):

There are 20 possible outcomes if I pull 2 marbles. red blue red green red orange red yellow blue red blue green and so on.

OpenStudy (anonymous):

Okay. And if you kept going it would be 3 then 2 then 1? So It would be like 5! right?

OpenStudy (anonymous):

Ah, yes! You've studied factorials =)

OpenStudy (anonymous):

So think of it this way. If I pull all 5, then it's 5! which is 5*4*3*2*1. But what happens if I don't pull all 5, but only 3 marbles? Well, then I don't want to do 5*4*3*2*1, I want to do just 5*4*3.

OpenStudy (anonymous):

Or if there were 10 marbles and I pulled 3, then it would be 10*9*8

OpenStudy (anonymous):

Yes I have studied them, I'm in the middle now actually. And if you didn't pull them all you would only multiply the first x number? x=number of marbles pulled. Right?

OpenStudy (anonymous):

Very good =) Exactly. Now, what I'm going to show you next is just a fancy way to write that as an equation, but you have the basic concept.

OpenStudy (anonymous):

Let's say I have 10 marbles and I pull 4. If I do 10!, that's too much. I need to stop before multiplying all of those. There's a simply way to write an equation that says WHEN to stop. \[\frac{10!}{6!} = \frac{10*9*8*7*6*5*4*3*2*1}{6*5*4*3*2*1}\]

OpenStudy (anonymous):

Okay, so I know that last step maybe looks just random and crazy, but look what happens when I divide 10! by 6!. Do you see how 6 is in the top and bottom? Do you see how everything after 6 is also in the top and bottom?

OpenStudy (anonymous):

Yes. And so it cancels out leaving only what you need. 10*9*8*7, right?

OpenStudy (anonymous):

Exactly right =) So when I wrote \[\frac{10!}{6!}\] it was a fancy, but neat way of writing 10 factorial, but stop at 7, don't do 6 or anything after.

OpenStudy (anonymous):

Now, let's try to answer their question Type the formula for the number of permutations of n things taken r at a time. P(n,r)=? What if there are n marbles in the bag, and we pull r of them?

OpenStudy (anonymous):

p=n!/r!?

OpenStudy (anonymous):

Would that work for our 10 marble bag when we pulled out 4? It would be 10!/4!, which would mean we would do 10*9*8*7*6*5

OpenStudy (anonymous):

Hmm. No, so.. P=n!/(n-r)! maybe..?

OpenStudy (anonymous):

That's the only way I could think to get 6 on the bottom.

OpenStudy (anonymous):

Oh my goodness. So impressive.

OpenStudy (anonymous):

Wait. Really? No way.

OpenStudy (anonymous):

Really. Has anyone ever told you that you're brilliant at math?

OpenStudy (anonymous):

No. There is probably a reason for that. Haha. And thank youuu!!!